What is the radius of Li 2+ ion in 2nd orbit and please give me the formula

Grade:11

9 Answers

Umakant biswal
7 years ago
@ piyush 
u haVE to use the below formula 
r (n ) = n ^ 2*0.529* 10^ -10 / zm 
z = atomic number = 3 
n = 2 (as given 2nd orbit ) 
just put the value in the formula and calculate it . 
r (n ) = 4 * 0.529* 10 ^ -10 / 3
and u will get the answer . 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
Lakshit
7 years ago
The formula of finding the radius of atom = 0.529 * n^2/Z
So, here the ‘n’ is 2 & ‘Z’ is 3
Then, Radius=0.529 * 4/3
                   =0.705
Nikhil Goyal
6 years ago
Formula For the radius is- n^2/Z×0•529×10^-10mIn this question,Z=3,n=2Using formula4/3×0•529×10^-10=0•7053×10^-10✓✓✓Best Of Luck✓✓✓
Rahul chouhan
6 years ago
formula of finding the radius of atom = 0.529 * n^2/ZSo, here the ‘n’ is 2 & ‘Z’ is 3Then, Radius=0.529 * 4/3 =0.705 Ans
Simran Singh
4 years ago
For finding the radius of 2nd excited state in doubly ionised lithium n=3 because 2nd excited state means 3 orbital &z=3 so        _13.6[3^2/3]=1.587A
ertagul
3 years ago
formula of finding the radius of atom = 0.529 * n^2/ZSo, here the ‘n’ is 2 & ‘Z’ is 3Then, Radius=0.529 * 4/3 =0.705 Ans by it self                                                                                                                                                                                      Answer is =0.705
 
ertagul
3 years ago
formula of finding the radius of atom = 0.529 * n^2/ZSo, here the ‘n’ is 2 & ‘Z’ is 3Then, Radius=0.529 * 4/3 =0.705 Ans
ertagul
3 years ago
@ piyush 
u haVE to use the below formula 
r (n ) = n ^ 2*0.529* 10^ -10 / zm 
z = atomic number = 3 
n = 2 (as given 2nd orbit ) 
just put the value in the formula and calculate it . 
r (n ) = 4 * 0.529* 10 ^ -10 / 3
and u will get the answer . 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST 
Yash Chourasiya
3 years ago
Dear Student

As we know,
rn​= r1 for H × n2/Z
= 0.529A×n2/Z​
r2​for Li2+= 0.529A×22/3​
= 0.529A×4​/3 A0

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya